# Anole: Adapting Diverse Compressed Models for Cross-scene Prediction on Mobile Devices

### v1

@@ Abstract
Index Terms:
Model inference, online algorithms, mobile devices, cross-scene, out of distribution, reliability
@@ I Introduction
Motivation. Last decade has witnessed the booming development of Artificial Intelligence of Things (AIoT), an emerging computing paradigm that marries artificial intelligence (AI) and Internet of Things (IoT) technologies to enable independent decision-making at each component level of the interconnected system. In many AIoT scenarios, deep neural network (DNN) model inference (i.e., prediction) tasks are required to execute on mobile devices, referred to as the online mobile inference (OMI) problem, with stringent accuracy and latency requirements. For example, unmanned aerial vehicles (UAVs) need to constantly detect surrounding objects in real time [1]; a dash cam mounted on a vehicle needs to perform continuous image object detection [2]; robots in smart factories need to detect objects in production lines in real time, and interact with human workers and other robots [3].
@@
To address the OMI problem, however, is demanding for two reasons as follows. First, given that mobile devices constantly experience scene changes while moving (e.g., due to various lighting conditions, weather conditions, and viewing angles), the output of DNNs should remain reliable and accurate. Training a statistical learning DNN on a given dataset, as in normal deep learning paradigm, becomes difficult to guarantee the robustness, interpretability and correctness of the output of the statistical learning models when data samples are out-of-distribution (OOD) [4]. Second, the response time for model inference should satisfy a rigid delay budget to support real-time interactions with these devices. As mobile devices are resource-constrained in terms of computation, storage and energy, they cannot handle large DNNs. Though it would be beneficial to offload a part of or even entire inference tasks to a remote cloud, unstable communication between mobile devices and the cloud may lead to unpredictable delay.
@@
In the literature, much effort has been made to improve DNN model inference accuracy on mobile devices but in static scenarios. One main branch aims to develop DNNs specially designed for mobile devices [5, 6, 7, 8] or to compress (e.g., via model pruning and quantization) existing DNNs to match the computing capability of a mobile device [9, 10]. Such schemes ensure real-time model inference at the expense of compromised accuracy, especially when dealing with OOD data samples. Another branch is to divide DNNs and perform collaborative inference on both edge devices and the cloud [11, 12, 13], or to transmit compressed sensory data to the cloud for data recovery and model inference [14, 15]. These approaches need coordination with the cloud for each inference, leading to unpredictable inference delays when the communication link is unstable or disconnected. As a result, to the best of our knowledge, there is no successful solution to the OMI problem yet.
@@
We implement Anole on three typical mobile devices, i.e., Jetson Nano, Jetson TX2 NX and a laptop, with each equipped with a CPU/MCU and an entry-level GPU, to conduct the image object detection task on moving vehicles. Specifically, we train the $\mathcal{M}_{scene}$ based on Resnet18 [16], a pack of 19 compressed DNNs based on YOLOv3-tiny [17], and the $\mathcal{M}_{decision}$ based on Resnet18 accordingly, using three driving video datasets collected from multiple cities in different counties. We conduct extensive trace-driven and real-world experiments using UAVs. Results demonstrate that Anole is lightweight and agile to switch best models with low latencies of 61.0 ms, 13.9 ms, and 52.0 ms on Jetson Nano, Jetson TX2 NX, and the laptop, respectively. In cross-scene (i.e., seen but fast-changing scenes) setting, Anole can achieve a high F1 prediction accuracy of 56.4% whereas the F1 score of a general large DNN model and a general compact DNN are 50.7% and 45.9%, respectively. In hard new-scene (i.e., unseen scenes) setting, Anole can maintain a high F1 score of 48.7% whereas the F1 score of the general large DNN and the general compact DNN drops to 46.6% and 41.1%, respectively.
We highlight the main contributions made in this paper as follows: 1) A new solution to the OMI problem by recruiting a pack of compact but specialized models on resource-constrained mobile devices, without any intervention with the cloud during online model inference; 2) A scene partition method that effectively facilitates the training of specialized models by leveraging both semantic and feature similarity of the data; 3) We have implemented Anole on typical mobile devices and conducted extensive trace-driven and real-world experiments, the results of which demonstrate the efficacy of Anole.
@@ II-B Problem Formulation
III Overview of Anole
The design of Anole is motivated by an observation that though any single compressed model generally has a lower prediction accuracy than the big model, there exists a compressed model that can achieve comparable accuracy as the big model for each specific scene. As illustrated in Figure 2, Anole consists of two parts, i.e., offline scene profiling and online model inference.
@@ III Overview of Anole
IV Offline Scene Profiling
IV-A Training Compressed Models
IV-A1 Training Dataset Segmentation
@@
IV-A2 Compressed Model Training
@@
Input: Semantic-defined scenes $\Gamma_{i}^{sem}$ for $i\in[1,m]$, preset number $n$ of compressed models to be trained, threshold $\delta$ at which the model performance meets the required criteria.
@@
// Scene embedding.
1 Train the scene embedding model $\mathcal{M}_{scene}$ with the supervision of semantic information defined in $\Gamma_{i}^{sem}$ for $i\in[1,m]$;
2 for each $\Gamma_{i}^{sem}$ do
3       $H_{i}\leftarrow\mathcal{M}_{scene}(\Gamma_{i}^{sem})$;
4
// Model training with multi-level clustering.
5 Compressed model repository $M^{rep}\leftarrow\{\}$, clustering number $k\leftarrow 2$;
@@ 16      $k\leftarrow k+1$;
Algorithm 1 Compressed Model Training Algorithm
@@
Model Training with Multi-level Clustering. Instead of training compressed models directly from $\Gamma_{i}$ for $i\in[1,m]$, we further consider the feature similarity of data samples by clustering embeddings in all $H_{i}$ and train compressed models on obtained clusters. Specifically, to obtain clusters with different levels of similarity, we conduct multiple $k$-means [18] clustering with $k$ varying from 2 over embeddings in all $H_{i}$ for $i\in[1,m]$. For each $k$, all embeddings can be divided into $k$ clusters, denoted as $H^{k}_{j}$ for $j\in[1,k]$. We train a compressed model, denoted as $\mathcal{M}^{k}_{j}$, on each clustered scene corresponding to $H^{k}_{j}$ for $j\in[1,k]$ and validate its performance. If the performance of $\mathcal{M}^{k}_{j}$ exceeds a threshold $\delta$, $\mathcal{M}^{k}_{j}$ is added to the compressed model repository. This procedure repeats until a set of $n$ compressed models $\{\mathcal{M}_{1},\mathcal{M}_{2},\cdots,\mathcal{M}_{n}\}$ are derived, where $n$ denotes a preset number for compressed models to be trained.
IV-B Adaptive Scene Sampling
@@
To obtain $\{\Psi_{1}^{sub},\Psi_{2}^{sub},\cdots,\Psi_{n}^{sub}\}$, a straightforward idea is to randomly pick a number of samples $X$ from $D$ and test $\mathcal{M}_{i}$ for $i\in[1,n]$. If a $\mathcal{M}_{i}$ can achieve satisfactory prediction accuracy on sample $x\in X$, $x$ belongs to $\Psi_{i}^{sub}$. As $\{\Psi_{1},\Psi_{2},\cdots,\Psi_{n}\}$ may be biased in $D$, such random sampling algorithm also generates unbalanced $\{\Psi_{1}^{sub},\Psi_{2}^{sub},\cdots,\Psi_{n}^{sub}\}$. To solve the unbalanced sampling problem, however, is not intuitive, because of Proposition 1. Proposition 1 holds that we can not know a sample belongs to which distribution from all the distributions those models can characterize (i.e., $\{\Psi_{1},\Psi_{2},\cdots,\Psi_{n}\}$) without high computational cost experiments. In order to obtain a balanced $\{\Psi_{1}^{sub},\Psi_{2}^{sub},\cdots,\Psi_{n}^{sub}\}$ at a low computation cost, we design an adaptive sampling algorithm based on Thompson sampling [19].
@@
Specifically, in the $k$-th sampling round for $k\in\mathbb{N}$, we first examine if the training set $\Gamma_{i}$ of $\mathcal{M}_{i}$ for $i\in[1,n]$ has been well sampled by checking
$|S_{i}|>\frac{\log(1-\theta^{\frac{1}{|\Gamma_{i}|}})}{\log(1-\frac{1}{|\Gamma_{i}|})},$
where $S_{i}$ is the set of samples sampled from $\Gamma_{i}$; $\theta$ is the confidence of being well sampled; and ||$\cdot$|| ⋅ | is the number of elements in a set.
@@
(a) Random sampling
(b) Adaptive sampling
IV-C Training Decision Model

### v2

# A Scene-aware Models Adaptation Scheme for Cross-scene Online Inference on Mobile Devices

@@ I Introduction
Motivation. Last decade has witnessed the booming development of Artificial Intelligence of Things (AIoT), an emerging computing paradigm that marries artificial intelligence (AI) and Internet of Things (IoT) technologies to enable independent decision-making at each component level of the interconnected system. In many AIoT scenarios, deep neural network (DNN) model inference (i.e., prediction) tasks are required to execute on mobile devices, referred to as the online mobile inference (OMI) problem, with stringent accuracy and latency requirements. For example, unmanned aerial vehicles (UAVs) need to constantly detect surrounding objects in real time [wang2023generalized]; a dash cam mounted on a vehicle needs to perform continuous image object detection [zhou2021monoef]; robots in smart factories need to detect objects in production lines in real time, interact with human workers and other robots [zheng2020contextual] or even make production decisions [dalmia2020pest].
@@
To address the OMI problem, however, is demanding for two reasons as follows. First, given that mobile devices constantly experience scene changes while moving (e.g., due to various lighting conditions, weather conditions, and viewing angles), the output of DNNs should remain reliable and accurate. Training a statistical learning DNN on a given dataset, as in normal deep learning paradigm, becomes difficult to guarantee the robustness, interpretability and correctness of the output of the statistical learning models when data samples are out-of-distribution (OOD) [chen2019edge]. Second, the response time for model inference should satisfy a rigid delay budget to support real-time interactions with these devices. As mobile devices are resource-constrained in terms of computation, storage and energy, they cannot handle large DNNs. Though it would be beneficial to offload a part of or even entire inference tasks to a remote cloud, unstable communication between mobile devices and the cloud may lead to unpredictable delay.
@@
In the literature, much effort has been made to improve DNN model inference accuracy on mobile devices but in static scenarios. One main branch aims to develop DNNs specially designed for mobile devices [howard2017mobilenets, zhang2018shufflenet, sanh2019distilbert, jiao2019tinybert] or to compress (e.g., via model pruning and quantization) existing DNNs to match the computing capability of a mobile device [han2015learning, wang2020pruning]. Such schemes ensure real-time model inference at the expense of compromised accuracy, especially when dealing with OOD data samples. Moreover, even without considering the limited resources of mobile devices, it is also difficult to find a satisfactory neural network capable of stable inference on OOD samples for some complex tasks such as mobile decision making [zheng2022leon, zheng2020contextual]. Another branch is to divide DNNs and perform collaborative inference on both edge devices and the cloud [hu2019dynamic, kang2017neurosurgeon, fang2019teamnet], or to transmit compressed sensory data to the cloud for data recovery and model inference [Liu2019, zhang2021elf]. These approaches need coordination with the cloud for each inference, leading to unpredictable inference delays when the communication link is unstable or disconnected. As a result, to the best of our knowledge, there is no successful solution to the OMI problem yet.
@@
We implement Anole on three typical mobile devices, i.e., Jetson Nano, Jetson TX2 NX and a laptop, with each equipped with a CPU/MCU and an entry-level GPU, to conduct the image object detection task on moving vehicles. Specifically, we train the $\mathcal{M}_{scene}$ based on Resnet18 [he2016deep], a pack of 19 compressed DNNs based on YOLOv3-tiny [adarsh2020yolo], and the $\mathcal{M}_{decision}$ based on Resnet18 accordingly, using three driving video datasets collected from multiple cities in different counties. We conduct extensive trace-driven and real-world experiments using UAVs. Results demonstrate that Anole is lightweight and agile to switch best models with low latencies of 61.0 ms, 13.9 ms, and 52.0 ms on Jetson Nano, Jetson TX2 NX, and the laptop, respectively. In cross-scene (i.e., seen but fast-changing scenes) setting, Anole can achieve a high F1 prediction accuracy of 56.4% whereas the F1 score of a general large DNN model and a general compact DNN are 50.7% and 45.9%, respectively. In hard new-scene (i.e., unseen scenes) setting, Anole can maintain a high F1 score of 48.7% whereas the F1 score of the general large DNN and the general compact DNN drops to 46.6% and 41.1%, respectively.
We highlight the main contributions made in this paper as follows:
  • A new solution to the OMI problem by recruiting a pack of compact but specialized models on resource-constrained mobile devices, without any intervention with the cloud during online model inference;
  • A scene partition method that effectively facilitates the training of specialized models by leveraging both semantic and feature similarity of the data;
  • We have implemented Anole on typical mobile devices and conducted extensive trace-driven and real-world experiments on 3 typical tasks, the results of which demonstrate the efficacy of Anole.
@@ II-B Problem Formulation
III Empirical Study
We first investigate how DNNs behave in mobile settings, taking the typical online object detection as the example task.
III-A Driving Video Datasets
We consider three representative driving video/image datasets as follows:
  • KITTI [geiger2012we]: comprises 389 stereo and optical flow image pairs, stereo visual odometry sequences of 39.2 km length, and more than 200k 3D object annotations captured in cluttered scenarios (up to 15 cars and 30 pedestrians are visible per image). For online object detection, KITTI consists of 21 training sequences and 29 test sequences.
  • BDD100k [bdd100k]: contains over 100k video clips regarding ten autonomous driving tasks. Clips of 720p and 30fps were collected from more than 50 thousand rides in New York city and San Francisco Bay Area, USA. Each clip lasts for 40 seconds and is associated with semantic attributes such as the scene type (e.g., city, streets, residential areas, and highways), weather condition and the time of the day.
  • SHD: contains 100 driving video clips of one minute recorded in March 2022 with a 1080p dashcam in Shanghai city, China. Clips were collected from ten typical scenarios, including highway, typical surface roads, and tunnels, at different time in the day. LabelImg [tzutalin2015labelimg] is employed to label objects in all images.
(a) CDF of image brightness
(b) CDF of image contrast
(c) CDF of the number of objects
(d) CDF of the ratio of object area
III-B Mobile Inference Accuracy Analysis
[…]
(a) Boxplot of F1 score, where $\mathcal{M}_{best}$ denotes the best compressed models of each scene
(b) Heatmap of normalized F1 score tested on ten selected scenes defined with semantic attributes
(c) Heatmap of normalized F1 score tested on ten selected scenes defined based on feature similarity
Figure 3: A best-fit compressed model can achieve comparable prediction accuracy provided by a large DNN, whereas compressed DNNs trained using existing scene partitioning methods fail to perform well on their corresponding training dataset.
Then, we further define ten distinct scenes, denoted as $\Gamma_{i}$ for $i\in[1,10]$, based on the associated semantic attributes of all available training samples, such as {good weather, daytime, urban}, and for each scene, we train a compressed object detection models, denoted as $\mathcal{M}_{i}$ for $i\in[1,10]$.
Figure 3(a) plots the boxplot of the F1 score obtained with different models on all the 58 testing sets of seen clips. It can be seen that the general compressed model $\mathcal{M}_{0}$ is not the best compressed model. Moreover, we have the following observation:
Observation 1
Though any single compressed model generally has a lower prediction accuracy than the big model, there exists a compressed model that can achieve comparable accuracy as the big model for each specific scene.
From the above observation, it is possible to achieve appealing inference accuracy of the big model at a low cost of a compressed model, if we can identify an appropriate compressed model that best fits the current test sample. An intuitive model selection scheme is to choose the compressed model $\mathcal{M}_{k}$ trained on the dataset $\Gamma_{k}$ with similar semantic attribute values to the current test sample. This idea assumes that $\Gamma_{k}\subset\Psi_{k}$. Figure 3(b) plots the normalized F1 score for each $\mathcal{M}_{k}$ tested on each $\Gamma_{k}$. It can be seen that the highest F1 scores do not consistently appear along the diagonal, which makes the intuitive model selection scheme hard to work. Another scheme is to define scenes based on feature similarity, where data samples with similar features are clustered. Given identified scenes, compressed DNNs can be trained and the best DNN can be selected according to the feature similarity between the test samples and existing scenes. Similar model selection results can be seen in Figure 3(c).
IV Overview of Anole
As illustrated in Figure 4, Anole consists of two parts, i.e., offline scene profiling and online model inference.
@@ IV Overview of Anole
V Offline Scene Profiling
V-A Training Compressed Models
V-A1 Training Dataset Segmentation
@@
V-A2 Compressed Model Training
@@
Model Training with Multi-level Clustering. Instead of training compressed models directly from $\Gamma_{i}$ for $i\in[1,m]$, we further consider the feature similarity of data samples by clustering embeddings in all $H_{i}$ and train compressed models on obtained clusters. Specifically, two versions of multi-level clustering algorithms are proposed in Anole, i.e., Compressed Model Training with Distinct Samples among Models (CMT-DM, described in Algorithm 1), and Compressed Model Training with Shared Samples among Models (CMT-SM, described in Algorithm 2).
Input: Semantic-defined scenes $\Gamma_{i}^{sem}$ for $i\in[1,m]$, preset number $n$ of compressed models to be trained, threshold $\delta$ at which the model performance meets the required criteria, embeddings of all scenes $H_{i}$ for $i\in[1,m]$.
@@
1
1ex // Model training with multi-level clustering.
2 Compressed model repository $M^{rep}\leftarrow\{\}$;
3 clustering number $k\leftarrow 2$;
@@ 14 $k\leftarrow k+1$;
Algorithm 1 CMT-DM Algorithm
@@
The core idea of the CMT-DM algorithm is to obtain the clusters with different levels of similarity. To this end, we conduct multiple $k$-means [lloyd1982least] clustering with $k$ varying from 2 over embeddings in all $H_{i}$ for $i\in[1,m]$. For each $k$, all embeddings can be divided into $k$ clusters, denoted as $C^{k}_{j}$ for $j\in[1,k]$ (Line 4). We train a compressed model, denoted as $\mathcal{M}^{k}_{j}$, on each clustered scene corresponding to $C^{k}_{j}$ for $j\in[1,k]$ (Line 5) and validate its performance. If the performance of $\mathcal{M}^{k}_{j}$ exceeds a threshold $\delta$, $\mathcal{M}^{k}_{j}$ is added to the compressed model repository (Line 6 - Line 10). This procedure repeats until a set of $n$ compressed models $\{\mathcal{M}_{1},\mathcal{M}_{2},\cdots,\mathcal{M}_{n}\}$ are derived, where $n$ denotes a preset number for compressed models to be trained.
Input: Semantic-defined scenes $\Gamma_{i}^{sem}$ for $i\in[1,m]$, preset number $n$ of compressed models to be trained, threshold $\delta$ at which the model performance meets the required criteria, embeddings of all scenes $H_{i}$ for $i\in[1,m]$, initial clustering number $K$.
1
1ex // Model training with multi-level clustering.
2 Compressed model repository $M^{rep}\leftarrow\{\}$;
3 $\bm{C}=\{C_{1},C_{2},\cdots,C_{K}\}\leftarrow$ clusters on $\{H_{1},H_{2},\cdots,H_{m}\}$ with clustering number $K$;
4 while $|M^{rep}|<n$ do
5   Train $K$ compressed models $\mathcal{M}_{j}$ for $j\in[1,K]$ on clusters $\{C_{1},C_{2},\cdots,C_{K}\}$;
 // performance evaluation for each candidate compressed models.
6 for each $\mathcal{M}^{k}_{j}$ do
8    if $p_{j}>\delta$ then
9       $\mathcal{M}_{|M^{rep}|+1}\leftarrow\mathcal{M}^{k}_{j}$;
10       $M^{rep}$.append($\mathcal{M}_{|M^{rep}|+1}$);
11
12
13 $\bm{C}^{s}=\{C_{1}^{s},C_{2}^{s},\cdots,C_{K}^{s}\}\leftarrow$ ascending sorted clusters according to their performance;
 // clusters combination.
14 for each $C_{k_{1}}^{s}$ for $k_{1}\in[1,K]$ do
15    for each $C_{k_{2}}^{s}$ for $k_{2}\in[2,K]$ do
16       $p_{k_{1},k_{2}}\leftarrow$ performance of the model $\mathcal{M}_{k_{1},k_{2}}$ trained on $C_{k_{1},k_{2}}^{s}=C_{k_{1}}^{s}\cup C_{k_{2}}^{s}$;
17       if $p_{k_{1},k_{2}}>p_{k_{1}}$ and $p_{k_{1},k_{2}}>p_{k_{2}}$ then
18          $\bm{C}\leftarrow\bm{C}\cup\{C_{k_{1},k_{2}}^{s}\}\setminus\{C_{k_{1}}^{s},C_{k_{2}}^{s}\}$;
19
20       else if $p_{k_{1},k_{2}}>p_{k_{1}}$ and $p_{k_{1},k_{2}}\leq p_{k_{2}}$ then
21          $\bm{C}\leftarrow\bm{C}\cup\{C_{k_{1},k_{2}}^{s}\}\setminus\{C_{k_{1}}^{s}\}$;
22
23       else if $p_{k_{1},k_{2}}\leq p_{k_{1}}$ and $p_{k_{1},k_{2}}>p_{k_{2}}$ then
24          $\bm{C}\leftarrow\bm{C}\cup\{C_{k_{1},k_{2}}^{s}\}\setminus\{C_{k_{2}}^{s}\}$;
25
26
27
28
return $M^{rep}=\{\mathcal{M}_{1},\mathcal{M}_{2},\cdots,\mathcal{M}_{n}\}$
Algorithm 2 CMT-SM Algorithm
The core idea of the CMT-SM algorithm is to first cluster scenes at a fine-grained level and then to try to combine the clusters with bad model training performance by pairwise combination trail. To this end, we first conduct $k$-means clustering with a preset clustering number $K$ over embeddings in all $H_{i}$ for $i\in[1,m]$. All embeddings can therefore be divided into $K$ clusters, denoted as $\bm{C}=\{C_{1},C_{2},\cdots,C_{K}\}$, to get a fine-grained partitioning (Line 2). We train a model for each clusters on $\bm{C}$ and the models whose performance exceeds the threshold $\delta$ will be added to the compressed model repository (Line 5 - Line 9). Then, we sort $\bm{C}=\{C_{1},C_{2},\cdots,C_{K}\}$ in ascending validated performance order to obtain $\bm{C}^{s}=\{C_{1}^{s},C_{2}^{s},\cdots,C_{K}^{s}\}$ for further combination of clusters (Line 10). Next, we start from the two clusters with the worst training performance, i.e., $C_{1}^{s}$ and $C_{2}^{s}$, attempting to combine them pairwise, and verify whether the training performance, denoted as $p_{k_{1},k_{2}}$ of combined cluster, denoted as $C_{k_{1},k_{2}}^{s}=C_{k_{1}}^{s}\cup C_{k_{2}}^{s}$, is higher than that of $C_{1}^{s}$ and $C_{2}^{s}$, denoted as $p_{k_{1}}$ and $p_{k_{2}}$, respectively (Line 11 - Line 19). If $p_{k_{1},k_{2}}$ is larger than both $p_{k_{1}}$ and $p_{k_{2}}$, $C_{1}^{s}$ and $C_{2}^{s}$ will be removed from the clusters set $\bm{C}$ and $C_{k_{1},k_{2}}^{s}$ will be added; If only $p_{k_{1},k_{2}}$ is larger than only one of $p_{k_{1}}$ or $p_{k_{2}}$, only the less one ($p_{k_{1}}$ or $p_{k_{2}}$ which is less) will be removed from $\bm{C}$ and $C_{k_{1},k_{2}}^{s}$ will be added. This procedure repeats until a set of $n$ compressed models are derived.
V-B Adaptive Scene Sampling
(a) Random sampling
(b) Adaptive sampling
@@ V-B Adaptive Scene Sampling
To obtain $\{\Psi_{1}^{sub},\Psi_{2}^{sub},\cdots,\Psi_{n}^{sub}\}$, a straightforward idea is to randomly pick a number of samples $X$ from $D$ and test $\mathcal{M}_{i}$ for $i\in[1,n]$. If a $\mathcal{M}_{i}$ can achieve satisfactory prediction accuracy on sample $\in X, xx$ belongs to $\Psi_{i}^{sub}$. As $\{\Psi_{1},\Psi_{2},\cdots,\Psi_{n}\}$ may be biased in $D$, such random sampling algorithm also generates unbalanced $\{\Psi_{1}^{sub},\Psi_{2}^{sub},\cdots,\Psi_{n}^{sub}\}$. To solve the unbalanced sampling problem, however, is not intuitive, because of Proposition 1. Proposition 1 holds that we can not know a sample belongs to which distribution from all the distributions those models can characterize (i.e., $\{\Psi_{1},\Psi_{2},\cdots,\Psi_{n}\}$) without high computational cost experiments. In order to obtain a balanced $\{\Psi_{1}^{sub},\Psi_{2}^{sub},\cdots,\Psi_{n}^{sub}\}$ at a low computation cost, we design an adaptive sampling algorithm based on Thompson sampling [thompson].
@@
Specifically, in the $k$-th sampling round for $k\in\mathbb{N}$, we first examine if the training set $\Gamma_{i}$ of $\mathcal{M}_{i}$ for $i\in[1,n]$ has been well sampled by checking $|S_{i}|>\frac{\log(1-\theta^{\frac{1}{|\Gamma_{i}|}})}{\log(1-\frac{1}{|\Gamma_{i}|})}$, where $S_{i}$ is the set of samples sampled from $\Gamma_{i}$; $\theta$ is the confidence of being well sampled; and $|\cdot|$ is the number of elements in a set.
@@
V-C Training Decision Model